OpenStudy (anonymous):

28+7/4r I need to see step by step

OpenStudy (anonymous):

is it 7/4r or 7/4 times r?

OpenStudy (anonymous):

7/4r

OpenStudy (anonymous):

ok so what is it equal to?

OpenStudy (anonymous):

all it says is solve 28 + 7/4r=-28,sorry I forgot the -28

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so multiply both sides by 4r

OpenStudy (anonymous):

you get 28(4r) + 7 = -28/4r

OpenStudy (anonymous):

now simplify and it becomes 112r + 7 = -7/r

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

now multiply both sides by 1/r

OpenStudy (anonymous):

you then get 112 + 7/r = -7

OpenStudy (anonymous):

ow subtract 112 from both sides and get 7/r = -119

OpenStudy (anonymous):

now multiply both sides by - 1/7

OpenStudy (anonymous):

you get r = -119 times (-1/7)

OpenStudy (anonymous):

simplify and you get r = 17

OpenStudy (anonymous):

is that good?

OpenStudy (anonymous):

yes,ty